OpenStudy (anonymous):

Is y=1+(1/x)+(1/x^2) and odd or even function?

OpenStudy (amistre64):

what does f(-a) equal?

OpenStudy (hba):

y=f(x)=1+(1/x)+(1/x^2) Check for f(-x)

OpenStudy (amistre64):

or if numbers are more your type does f(-2) = f(2)? or does f(-2) = -f(2)??

OpenStudy (hba):

If @amistre64 1st condition gets proved its Even,If the 2nd condition gets proved its Odd,If your question does not satisfy the above conditions,Then its neither even nor odd.

OpenStudy (anonymous):

I got it. Thanks.

OpenStudy (amistre64):

or by analysis \[1+(1/x)+(1/x^2)=\frac{x^2+x+1}{x^2}\] these work like +- rules; even/even=even i believe

OpenStudy (amistre64):

but then again, the top is not even since it is shifted from the y axis

OpenStudy (hba):

Your'e welcome.

OpenStudy (hba):

In some case if the power turns out to be even then the function is even,if the power is odd then the answer is odd. The trigonometric values also : sin (–x) = –sin x cos (–x) = cos x tan (–x) = –tan x csc (–x) = –csc x sec (–x) = sec x cot (–x) = –cot x